The solution set of the equation x(x)-4x-12=0
Answer by nerdybill(7384)

x(x)-4x-12=0
or
x^2-4x-12=0
is a "quadratic" --
Two ways to solve:
- factoring (try this first)
- quadratic equation (last resort)
.
But your equation can be factored as:
(x-6)(x+2) = 0
Setting each of the terms inside the parenthesis to zero gives your solutions as:
x = {-2, 6}